Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 173     Entry time: 04 May 2023, 06:41
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 12.4 keV 200 pulses / train, relatively stable intensity ~0.8 mJ / pulse
  • took some time in changing X-ray energy (from 19.16 keV to 12.4 keV)

Optical laser delivery

  • 420 nm PP laser
  • ns laser for jet imaging

Achievements/Observations

  • M1 & M2 offset mirror optimized
  • NKB focus roughly optimized (we used defocused X-ray beam, about 2um)
  • Collected several runs for inside nozzle pump excitation project (both adaptive gain and fixed MG modes)
  • Nozzle damage with PP laser tested (before and after the night shift)

Issues

  • AGIPD FPGA (2nd hemisphere) issue (requied a power cycle)
ELOG V3.1.4-7c3fd00